OpenStudy (anonymous):

sqrt(3) divided by sqrt (12)-sqrt(3)=??

OpenStudy (lgbasallote):

hint \[\sqrt {12} = \sqrt 3 \times \sqrt 4\] does that give you any ideas @nhernandez6 ??

OpenStudy (anonymous):

I know that 3*4 would be 12 but im not sure on how to work it out

OpenStudy (btaylor):

you have:\[\sqrt{3}/(\sqrt{12}-\sqrt{3})\] multiply by \[(\sqrt{12}+\sqrt{3})/(\sqrt{12}+\sqrt{3})\]The denominator should cancel out, and you get 15 as your new denominator. For the numerator, you get \[\sqrt3(\sqrt{12}+\sqrt{3}) = \sqrt{36} + 3 = 6 + 3 = 9\] and 9/15 = 3/5

OpenStudy (anonymous):

so 3/5 would be the final answer?

OpenStudy (btaylor):

Yep.

OpenStudy (anonymous):

one question how did u get the 15?

OpenStudy (anonymous):

Thanks a bunch

OpenStudy (ajprincess):

\[(\sqrt{12}-\sqrt{3})(\sqrt{12}+\sqrt{3})=(\sqrt{12})^{2}-(\sqrt{3})^{2}=12-3=9\]

OpenStudy (btaylor):

Oops, my bad.

OpenStudy (ajprincess):

so 9/9=1

OpenStudy (anonymous):

\[\frac{\sqrt{ 3 }}{ \sqrt{ 12 }-\sqrt{ 3 } }\\\\= \frac{\sqrt{ 3 }}{ \sqrt{4*3}-\sqrt{ 3 } }\\\\= \frac{\sqrt{ 3 }}{ 2\sqrt{ 3 }-\sqrt{ 3 } }\\\\= \frac{\sqrt{ 3 }}{ \sqrt{ 3} }\\=1 \]